OpenStudy (anonymous):

These 2 questions go hand in hand with each other 3x2 + 52 = 4 Step 2 to solve the equation would be: Select one: a. Multiply each side by 3 ((Incorrect)) b. Divide each side by 3 c. Take the multiplicative inverse of 4 d. Add 52 to each side The final answer using imaginary and complex numbers would be: Select one: a. x = 16i Incorrect b. x = 4i c. x = ±4i d. No solution

OpenStudy (surana):

Is that 3 x 2 or 3x2?

OpenStudy (anonymous):

3 multiplied by 2?

OpenStudy (surana):

Wait a minute, is the 2 by any chance an exponent?

OpenStudy (anonymous):

Oh yes! sorry the 2 is an exponent sorry

OpenStudy (surana):

Yeah, I was confused there.

OpenStudy (surana):

Okay, I think you're supposed to get the variable (x) isolated. So theoretically, you have to take 52, remove that and throw it on the other side to make a -52.

OpenStudy (anonymous):

i know you subtract 52 on both sides but whats the next step i thought it was multiply 3 on each side

OpenStudy (surana):

I'm pretty sure it's divide, since I remember something along the lines of "make exponent as small as possible."

OpenStudy (anonymous):

oh okay and if you divide what do you get?

OpenStudy (surana):

Divide -48 by 3.

OpenStudy (anonymous):

its -16 i put +16 but thats wrong so itd be D no solution

OpenStudy (surana):

And I think we've just solved your problem.

OpenStudy (anonymous):

can you help me with this question.. An imaginary number: Select one: a. Is a real number Incorrect b. Is a real solution c. Is undefined d. None of the above

OpenStudy (surana):

Sure.

OpenStudy (anonymous):

thanks !

OpenStudy (surana):

https://en.wikipedia.org/wiki/Imaginary_number Imaginary numbers are like, these guys.

OpenStudy (anonymous):

so D?

OpenStudy (surana):

Possibly.

jimthompson5910 (jim_thompson5910):

`its -16 i put +16 but thats wrong so itd be D no solution ` be careful @LegendaryNikki just because there are no real solutions does not mean there are no solutions. An imaginary/complex solution is a solution

OpenStudy (anonymous):

possibly?

OpenStudy (surana):

I'm not entirely too sure.

OpenStudy (anonymous):

yeah i used that defintion to answer it i got it wrong but can you help me with this one realll quick? 10x^2 = –90 i got -9 and it was wrong..

jimthompson5910 (jim_thompson5910):

you forgot to take the square root of both sides

OpenStudy (surana):

Are there other answers available?

OpenStudy (anonymous):

a. ±3i b. 3 c. –3i

OpenStudy (surana):

Isn't the square root of ten like 3 point something?

OpenStudy (anonymous):

3.16

OpenStudy (surana):

And now I am genuinely confused.

OpenStudy (surana):

@Nnesha , would it be possible for you to help?

OpenStudy (anonymous):

me too

Nnesha (nnesha):

jim_thompson5910 -->> `you forgot to take the square root of both sides` so ye...... take square root both sides to cancel out the square suppose if it was x^3 then we have to take cube root both sides to cancel out the 3

OpenStudy (anonymous):

im still lost so its 3?

Nnesha (nnesha):

do you know about imaginary number ?? i = what ?

OpenStudy (surana):

So 3i means 3-what?

Nnesha (nnesha):

ohh well i was expecting an answer from here but ye use the fact \[\rm i=\sqrt{-1}\]

Nnesha (nnesha):

her*

OpenStudy (anonymous):

-3i

OpenStudy (surana):

That's the answer?

Nnesha (nnesha):

here is an example \[\large\rm y^2 \rightarrow \sqrt{y^2} \rightarrow y^{2*\frac{ 1 }{ 2 }}\] we can convert radical to an exponent that's how when we take square root it cancels out the square \[\large\rm y^{\cancel{2}*\frac{ 1 }{ \cancel{2} }}= y\] another example \[\sqrt{y^2}= 4 \rightarrow y= \pm2\]

Nnesha (nnesha):

how did you get -3i ?

OpenStudy (anonymous):

Thank you guys !!

OpenStudy (anonymous):

im sorry i gotta do something else but thanks so much ill figure it out in a little !!

OpenStudy (surana):

Wait, -3i is the answer?

Nnesha (nnesha):

\[\rm \sqrt{-9} ~can~be~written~as~\sqrt{-1}*\sqrt{9}\] use the fact i =sqrt{-1} i guess u can take it from hre

OpenStudy (surana):

I have absolutely no idea what I"m doing with this particular problem, actually.

OpenStudy (zarkon):

\[10x^2 = –90 \Rightarrow x^2=-9\Rightarrow x=\pm\sqrt{-9}\Rightarrow x=\pm3i\]

Nnesha (nnesha):

there we go.

OpenStudy (surana):

And problem solved.

OpenStudy (anonymous):

what about that other one were we got -16? i still dont understand that one

OpenStudy (zarkon):

it\s done the same way i did the one above